agnela22:

can you pick up a solid steel ball of radius 6 inches? steel has a density of 0.28 pounds per cubic inch. To the nearest pound, what is the weight of the ball?

surjithayer:

first find volume of sphere. \[volume~of~sphere=\frac{4}{3} \pi r^3\] mass of spherical ball =density *volume.

surjithayer:

in case it is a cylindrical ball of radius 6 inches and height 6 inches,then volume\[=\pi r^2 h\]